# “Confidently Nonsensical?”: A Critical Survey on the Perspectives and Challenges of ‘Hallucinations’ in NLP

### 11 Apr 2024 v1

@@ 5 Practitioner Survey of Hallucination
In this section, adopting a community-centric approach Narayanan Venkit (2023), we conduct a survey to gain insights into researchers’ perceptions on hallucinations in NLP. We aim to understand how frequently researchers utilize these models, how often they encounter hallucinations, and their firsthand experiences with model hallucinations.
The primary goal is to demonstrate how researchers and practitioners within the field perceive the concept of ‘hallucination’ and to expand our findings beyond the limitations of existing literature. This survey-based approach allows us to delve into the practical aspects of hallucination research and gather real-world perspectives from individuals actively engaged in NLP and AI research.
@@ 5.1 Survey Recruitment and Data Collection
To ensure a comprehensive view, we specifically targeted researchers familiar with AI and ML, primarily from disciplines such as computer science and information science. However, we also welcomed participants from other domains to explore their perceptions of whether they had the literature understanding of the concept of hallucination. The survey was examined and approved by the Institutional Review Board (IRB) for ethical practices.
For the survey, we employed a systematic approach by randomly selecting 15 universities from the top 100 in the USA as per the 2023 US News and World Report rankings News (2023), to then reach out to potential participants from the field of NLP and AI, including faculty members and graduate students, to provide their insights through the survey. Additionally, we expanded our outreach efforts by leveraging social media platforms to reach out to practitioners within the networks of all the authors involved in the survey. We received a total of 221 responses, out of which 171 were complete and usable for analysis.
@@ 5.2 Survey Structure
The survey employed a combination of 12 open-ended and close-ended questions. For the analysis of open-ended questions, we utilized thematic analysis, drawing from established methodologies outlined in Blandford et al. (2016); Terry et al. (2017). The analysis involves identifying patterns, themes, and categories within the free-text responses provided by participants allowing us to uncover nuanced perspectives and key themes related to the research questions. The close-ended questions were analyzed using descriptive statistics to summarize and analyze the numerical data obtained from respondents. Throughout the analysis process, the research team made collective decisions regarding the retention, removal, or reorganization of themes derived from open-ended responses.
@@ 5.3 Survey Findings
Here, we summarize insights from our 171 survey responses to explore various perspectives on hallucinations in LLMs, including perceptions, weaknesses, and preferences. The breakdown of responses indicates that 76.54% of participants were from academia, 20.98% from the industry, and 2.47% selected both.
Participants were also asked about their research area’s direct relation to AI and NLP. The analysis revealed that more than 68.52% of researchers indicated that their work is directly related to AI, while the remaining respondents either exhibited familiarity with or indirectly incorporated NLP and AI methodologies in their work. This highlights the substantial involvement of AI experts and practitioners in addressing issues related to LLM model hallucinations.
5.3.1 Weaknesses of LLM
Before delving into inquiries about hallucinations in LLMs, it is crucial to gain insights into the perceived weaknesses of these models from the participants’ perspective, as well as understand how frequently they utilize these models in their work.
The survey results indicate that a significant portion of researchers heavily utilize LLMs in their daily life. Specifically, 67.28% of respondents reported using these models atleast once a day, while 20.37% mentioned using them all the time, highlighting the ubiquity of these models.
Upon analyzing the themes derived from participants’ responses on the weaknesses of generative AI tools, it was observed that a substantial majority (55%) of researchers perceive the main weaknesses to be the generation of misinformation and hallucinations, despite both phenomena being essentially similar in nature. For instance,
[…]
I have been exploring these models to see what they get right and wrong. They get a lot of things wrong – what some people call “hallucinations”.—Emeritus Professor, NLP
Some of the other important weaknesses mentioned by the respondents are: biases, not following the prompts correctly, complex language, and not having a long memory. For example,
[…]
They produce a lot of inaccurate replies with great confidence. These models also tend to be very biased toward many socio-demographic groups.—Graduate student, GenerativeAI
[…]
It is hard to distinguish whether the information provided by them is accurate or not. Sometimes, the models generate text with reasoning making it sound convincing enough to be true - but ends up being incorrect ultimately.—Industry, GenerativeAI
The responses highlight a critical concern within the research community regarding the reliability and accuracy of outputs generated by LLMs, with potential implications for various applications and domains, providing us with a strong motivation behind this study.
The widespread use of LLMs, particularly prominent models such as GPT 3, 3.5, and 4, highlights their importance and impact on research and industry practices. However, it’s noteworthy that respondents also mentioned other LLM models that they use or are familiar with. These include Mistral, BERT, LLaMA2, Midjourney, ClaudeAI, Gemini, Vicuna, t5, Falcon, PaLM, Imagen, Dolly, Perplexity, among others.
@@ 5.3.2 Familiarity with Hallucination
The survey included the question on participants’ familiarity with the concept of ‘hallucinations’ in AI-generated text, measured on a 5-point Likert scale. The analysis revealed that 24.07% of researchers reported being extremely familiar with the concept, while 33.33% indicated being very familiar with it (Figure 3). Participants who indicated not being familiar with the term ‘hallucination’ (7.96%) also demonstrated implicit concerns with this phenomenon by highlighting issues such as generating incorrect responses, making mistakes, and crafting stories autonomously.
@@
5.3.3 Frequency of Encountering Hallucination
The survey included a question regarding the frequency of encountering ‘hallucinated’ content, defined as content that is factually incorrect or unrelated to the input, assessed on a 5-point Likert scale ranging from ‘Never’ to ‘Very frequently’ (Figure 4). The analysis revealed that 46.91% of respondents reported encountering hallucinated content occasionally, while 29.01% indicated experiencing it frequently. The results suggest that a substantial portion of researchers and practitioners encounter instances of hallucinated content in AI-generated outputs, indicating a prevalent issue in generative NLP models.
Figure 4: Frequency of Encountering Hallucination
@@ 5.3.5 Alternative Terms for Hallucination
This indicates that while the majority of respondents did not propose an alternative term, a notable proportion sees fabrications as a more suitable descriptor for the phenomenon of hallucination in AI-generated content. For example
@@ 5.3.6 Creativity and Positive Applications
5.3.7 Social Ramifications of Hallucination
Participants were prompted to explain the effects of hallucination on their work/daily life. The resulting themes, from our qualitative analysis of their inputs, are outlined below:
Not Good for Education: Respondents raised concerns about the extensive use of these models by students for homework, indicating potential negative impacts on their performance and learning abilities. The respondents believe that such reliance on these models can lead to a degradation in students’ learning. Additionally, respondents express skepticism about the suitability of these models for checking homework assignments.
[…]
I don’t actually use AI for my work; I just want to be aware of what it can do because my students are probably using it for their homework. It could have an impact on students’ mastery of the material.—Associate Prof, Biotechnology
Not Good for Scholarly Work: Several respondents noted that these models are not effective for scholarly purposes, citing instances where the models generated information that was not present in the original paper. They express concerns that if researchers rely on these models for tasks like literature summarizing, it could lead to a deterioration in scholarly processes. For example:
[…]
They tend to generate a lot of misinformed facts about certain groups or cultures that I have seen happen often. They also generate ’facts’ from scholarly works where the papers would not have mentioned the same.—Graduate student, NLP
Struggle with Code Generation: The models were deemed inefficient for code generation by multiple respondents, often producing code that lacks utility due to hallucinations. Respondents highlighted mismatches between the generated code and its intended purpose, emphasizing the need for thorough review before utilization. Various concerns were raised, including the loss of context during prolonged interactions, inaccuracies in complex coding tasks leading to erroneous outputs, fabrication of functions or attributes, inaccuracies in both code and associated theoretical concepts, necessitating extensive debugging and corrections, and a tendency to cycle back to previously incorrect suggestions despite error notifications.
[…]
I was asking an AI to generate me a piece of code. It ended up picking some code from one website and some from another and combining it. However those two websites (they were linked by chatgpt) we’re using different versions of the library so the resulting code couldn’t be executed.—Industry, Network and Security
Increase in Time for Task: A common sentiment among respondents is that these models frequently produce errors or false information, resulting in potential time wastage. While they acknowledge occasional helpfulness, there’s a consensus that reliance on these models can often lead to unfavorable outcomes, particularly when verifying outputs. This dependency on verification contributes to increased task duration, adding extra work and time toward the project’s conclusion, as noted by several respondents.
[…]
I use GPT API to conduct analysis for some of my work and accuracy and consistency would be good in my context, and I have to find ways to finetune it before I can trust the results of the analysis, which added more work on my end.—Graduate Student, HCI
Misleading and Distrust: Generating incorrect outputs with confidence can lead to the dissemination of non-existent knowledge, such as misleading information in the literature that may confuse individuals with incorrect concepts. Most of our respondents mentioned this concern. Moreover, it poses challenges in differentiating between accurate AI responses and hallucinations, particularly for users lacking expertise in the relevant subject matter.
[…]
It leads to problems if even I do not have any idea about the work. It is hard to differentiate if it is a genuine output or hallucination.—Graduate Student, Data Science
5.4 An External Viewpoint
Additionally, our survey of 51 researchers who do not specialize in AI revealed that all except 3 have used text-generation models like various versions of ChatGPT. Despite their fields not being directly related to AI, a significant number integrate these tools into their workflow, with 19.6% using them multiple times daily and 11.76% using them several times per hour. Their extensive usage has allowed them to identify several limitations in the models; they are: mathematical inaccuracy, outdated information, misinformation, poor performance with complex tasks and creative thinking, lack of specificity in-depth, overconfidence, lack of transparency, bias, and irrelevant responses.
Based on the definitions provided, it is observed that there is a lack of clarity among the respondents regarding what constitutes a ‘hallucination’ in generative AI models, with perspectives varying widely. Thematic analysis of their responses indicates that the predominant view associates ’hallucination’ with the generation of nonfactual content and misinformation by AI systems. That means these models are generating facts that are not real and misleading. The remaining themes are factually incorrect, biased outputs, incompleteness, misinformation with confidence, and nonsensical but good-looking texts.
The results demonstrate the unclear comprehension and significance attributed to hallucination in language models beyond the field of NLP and AI. There is a pressing need to enhance public understanding of the concept of hallucination, emphasizing its meaning and strategies for mitigation. Given the increasing prominence of language models as sociotechnical systems Narayanan Venkit (2023), it is crucial to grasp their social interactions and potential societal ramifications.
5.5 Additional Impacts and Concerns
We analyzed perceptions when participants were asked about any additional concerns during the survey. Participants emphasized the necessity for greater control and more nuanced mechanisms to address and manage AI hallucinations effectively. Presently, the detection and rectification of hallucinations rely heavily on meticulous human review, highlighting the need for tools designed specifically to identify and mitigate such occurrences. The presence of hallucinations can significantly impact the credibility and acceptance of generative models among the general public. These issues arise due to the inherent limitations of generative algorithms and the absence of access to real-time external knowledge.
Transparency regarding the limitations of generative AI is deemed essential through our findings, and user education is seen as a key factor in mitigating risks associated with the unchecked use of AI-generated content, as the responsibility for identifying hallucinations often falls on the user. While inaccuracies in non-critical applications, like movie suggestions, may be tolerable, according to our survey, they are deemed crucially problematic in contexts such as business decision-making, law, or health Dahl et al. (2024).
6 Research Challenges and Recommendations
@@
Finally, based on our audit and survey analysis, we outline the key weaknesses encountered in hallucination within NLP and potential recommendations motivated by the weaknesses. Our focus is to utilize a community-centric and literature-based approach to define the primary weaknesses of the field currently and a path forward.
@@ 6.1 Challenges
Wide range of vague and absent definitions: The literature and the practitioner’s survey show diverse and conflicting definitions and frameworks, often lacking clarity or omitting explicit definitions for hallucination and how it is perceived in various fields of NLP.
Lack of standardization in measurement: The absence of standardized metrics to quantify hallucination results in the use of multiple scales and categorizations. This lack of standardization makes it challenging to compare and interpret results across different models and studies, leading to a proliferation of diverse approaches for evaluating hallucinations.
Limited awareness of hallucination in a sociotechnical context: Hallucination research often lacks the understanding of how the concept of hallucination is conceptualized beyond its technical purview (discussed in Section 4). When such analysis is employed in sociotechnical systems like healthcare and policy making Dahl et al. (2024); Pal et al. (2023), it is important to define the socially relevant framework of hallucination as well. There is no motivation shown to understand the social, political, and psychological considerations of hallucination in these works.
Lack of consensus between various subfields There are multiple frameworks of hallucination created across various subfields in NLP. For instance, hallucination in image captioning exhibits distinct characteristics compared to abstract summarization and conversational AI. These frameworks have been adopted based on individual usage without reaching an accord among other existing works.
Multiple sentiment towards Hallucination The perception of hallucination in generative AI varies depending on the context. For instance, it is often positively regarded as creativity in image generation, whereas in text generation, it is viewed negatively as errors or mistakes. Consequently, future research efforts should aim to better address this disparity to develop a more nuanced framework for understanding hallucination.
@@
Ethical and Social Challenges: Given that researchers and users often rely on these models for critical decision-making processes, there is a need for well-defined policies regarding the extent of their utilization and the transparency surrounding their use and the impact of hallucinations within such models. Currently, such policies are lacking in industries and academic settings.
@@ 6.2 Recommendations
Prior works like Blodgett et al. (2020) & Venkit et al. (2023) have formulated thematic recommendations to address AI issues. Expanding on this, we examine strategies for NLP practitioners studying ‘hallucination’ to overcome these challenges. We propose two overarching themes with four associated recommendations.
6.2.1 Author-Centric Recommendation
@@
6.2.2 Community-Centric Recommendation
@@ 7 Conclusion
Our work delves into the conceptualization of hallucination within the scope of NLP. Our approach involved two key methodologies: first, an exhaustive audit of 103 peer-reviewed papers in the NLP domain, and second, a practitioner survey of 171 researchers and academics to gauge the field’s perception and understanding of hallucination. Through this analysis, we have gained insights into how the NLP community conceptualizes and defines hallucination. Additionally, our thematic and community-based approach highlights potential weaknesses within the field, particularly in addressing misrepresentations and inaccurate characterizations associated with hallucination. Our work finally contributes to a deeper understanding of the challenges and gaps in research related to hallucination within NLP, paving the way for future advancements in this area.
@@ 9 Ethics Statement
Data Publication: All the papers used in our research are listed in the Appendix. However, we recognize the importance of transparency and accountability. Therefore, we publish the complete collection of papers along with our qualitative classification and annotation, allowing for public scrutiny and examination 333github.com/PranavNV/The-Thing-Called-Hallucination.
@@
Appendix A Appendix
A.1 Related Surveys on Hallucination
This section provides an overview of several key surveys in the realm of NLP focusing on the topic of hallucination and why our work addresses a relevant gap in the field. Starting with Ji et al. (2023a), this survey extensively delves into the advancements and challenges concerning hallucination in NLG, distinguishing between intrinsic and extrinsic frameworks of hallucination. Additionally, it sheds light on fundamental terms such as hallucination, faithfulness, and factuality, along with prevalent metrics for quantifying these phenomena. Rawte et al. (2023b) categorizes existing works within the domain of language models, covering various aspects including methods for detecting hallucination, mitigation techniques, datasets used, and evaluation metrics. Zhang et al. (2023c) addresses the challenges of hallucination in LLMs by categorizing hallucinations into input-conflicting, context-conflicting, and fact-conflicting types, diverging from traditional viewpoints.
While these surveys offer insights into the current state of hallucination research, they often do not pay attention to critical examinations of the field’s weaknesses arising from a lack of discourse in defining hallucination. Even within these surveys, a noticeable lack of discourse is observed regarding the characterization of hallucination. This deficiency in discussion reflects the broader trend within the entire field.
A.2 Works and Application
We illustrate the examples and categories of works that were looked into for understanding the various applications of hallucinations. We categorize the research on hallucinations into 7 major categories. The definitions and categories of all the applications are mentioned in Table 1.
Abstractive Summarization: Zhang et al. (2019); Son et al. (2022); Maynez et al. (2020); Choubey et al. (2023); Cao et al. (2021); Marfurt and Henderson (2022); Akani et al. (2023); van der Poel et al. (2022); Chen et al. (2023b); Dong et al. (2022); Shen et al. (2023); Nan et al. (2021); Chen et al. (2021); Ladhak et al. (2023); Nan et al. (2021); Flores and Cohan (2024)
Conversational AI: Liu et al. (2022); Zhou et al. (2020); Ji et al. (2023b); Zhang et al. (2023b); Yang et al. (2023); Das et al. (2022); Bouyamourn (2023); Sun et al. (2023); Sadat et al. (2023); Slobodkin et al. (2023); Ramakrishna et al. (2023); Xiao and Wang (2021b); Shuster et al. (2021); Nie et al. (2019); Longpre et al. (2021); Dziri et al. (2022); Maheshwari et al. (2023); Ladhak et al. (2022); Xu et al. (2023); Chen et al. (2023a); Goldberg et al. (2022); Sundar and Heck (2023); Roller et al. (2021); Mielke et al. (2022); Roller et al. (2021); Massarelli et al. (2020); Weller et al. (2024); Smith et al. (2022)
Data Augmentation: Jian et al. (2022); Ji et al. (2023b); Friedl et al. (2021); Samir and Silfverberg (2022); Anastasopoulos and Neubig (2019); Narayanan Venkit et al. (2023)
Image and Video Captioning: Xiao and Wang (2021b); Dai et al. (2023); Rohrbach et al. (2018); Li et al. (2023); Testoni and Bernardi (2021); Son et al. (2022); Dai et al. (2023); Li et al. (2023); Liu and Wan (2023)
Machine Translation: Wang and Sennrich (2020); Raunak et al. (2021); Dale et al. (2022); Guerreiro et al. (2023a); Xu et al. (2023); Pfeiffer et al. (2023); Guerreiro et al. (2023b); Dale et al. (2023); Irvine and Callison-Burch (2014); Ferrando et al. (2022); Vu et al. (2022); Müller et al. (2020); Waldendorf et al. (2024)
Data2Text Generation: González Corbelle et al. (2022); Shi et al. (2023); Yoon et al. (2022); Filippova (2020b); Kothyari et al. (2023); Lango and Dusek (2023); Cirik et al. (2022); Fei et al. (2023); Obaid ul Islam et al. (2023); Qiu et al. (2023); Testoni and Bernardi (2021); González-Corbelle et al. (2022); Islam et al. (2023); Polat et al. (2023)
Miscellaneous: Manakul et al. (2023); Ji et al. (2023d); Maharaj et al. (2023); McKenna et al. (2023); Pal et al. (2023); Zhao et al. (2023); Berbatova and Salambashev (2023); Wu et al. (2024)
A.3 Supplementary Survey Analysis
Figure 5: Frequency of Text Generation Model Usage

### 14 Sep 2024 v2

@@ 5 Practitioner Survey of Hallucination
In this section, adopting a ‘community-centric approach’ Narayanan Venkit (2023), we conduct a survey to gain insights into researchers’ perceptions of hallucinations in NLP to complement our theoretical discussions with practical real-world perspectives. The primary goal is to demonstrate how researchers and practitioners within the field perceive the concept of ‘hallucination’ and to expand our findings beyond the limitations of existing literature where real-world perceptions from the researchers are missing Huang et al. (2023); Zhang et al. (2023c); Ji et al. (2023a). This motivates us to gather real-world perspectives from individuals actively engaged in NLP and AI research.
@@ 5.1 Survey Recruitment and Data Collection
To ensure a comprehensive view, we specifically targeted researchers familiar with AI and ML, primarily from disciplines such as computer science and information science. However, we also welcomed participants from other domains to explore their perceptions of whether they had the literature understanding of the concept of hallucination as they are also extensively using LLM models. The survey was examined and approved by the Institutional Review Board (IRB) for ethical practices.
We additionally employed a systematic approach by randomly selecting 15 universities from the top 100 in the USA as per the 2023 US News and World Report rankings News (2023), to then reach out to potential participants. Prior works Chakravorti et al. (2023) have previously employed this process to identify high-quality participants. We received a total of 223 responses, out of which 171 were complete and usable for analysis.
@@ 5.2 Survey Structure
The survey employed a combination of 14 open-ended and close-ended questions. The survey has been built based on the previous survey design techniques Rosen et al. (2013); Baker (2016); Van Noorden and Perkel (2023); Chakravorti et al. (2024). Open-ended questions and free-response text boxes allow us to gather rich opinions from participants. This approach integrates all our findings, providing a broader and deeper understanding of the response. For the analysis of open-ended questions, we utilized thematic analysis, drawing from established methodologies outlined in Blandford et al. (2016); Terry et al. (2017). The close-ended questions were analyzed using descriptive statistics to summarize and analyze the numerical data obtained from respondents. Throughout the analysis process, the research team made collective decisions regarding the retention, removal, or reorganization of themes derived from open-ended responses. All the survey questions have been provided in Appendix 10.4.
@@ 5.3 Survey Findings
We now summarize insights from our responses to explore various perspectives on hallucinations in LLMs, including perceptions, weaknesses, and preferences. The breakdown of responses indicates that 76.54% of participants were from academia, 20.98% from the industry, and 2.47% both.
Participants were also asked about their research area’s direct relation to AI and NLP. The analysis revealed that more than 68.52% of researchers indicated that their work is directly related to NLP, while the remaining respondents either exhibited familiarity with or indirectly incorporated NLP and AI methodologies in their work. This highlights the substantial involvement of AI experts and practitioners within the survey.
@@ 5.3.1 Familiarity with Hallucination
The survey included the question on participants’ familiarity with the concept of ‘hallucinations’ in AI-generated text, measured on a 5-point Likert scale. The analysis revealed that 24.07% of researchers reported being extremely familiar with the concept, while 33.33% indicated being very familiar with it (Figure 3). Participants who indicated not being familiar with the term ‘hallucination’ (7.96%) also demonstrated implicit concerns with this phenomenon by highlighting issues such as generating incorrect responses and crafting stories autonomously. This demonstrates the widespread impact of the phenomenon within the community.
@@
Figure 4: Frequency of encountering ‘Hallucination’
5.3.2 Hallucination Frequency
The survey included a question regarding the frequency of encountering ‘hallucinated’ content, defined as content that is factually incorrect or unrelated to the input, assessed on a 5-point Likert scale ranging from ‘Never’ to ‘Very frequently’ (Figure 4). The analysis revealed that 46.91% of respondents reported encountering hallucinated content occasionally, while 29.01% indicated experiencing it frequently. The results suggest that a substantial portion of practitioners encounter instances of hallucinated content in AI-generated outputs, indicating a prevalent issue in generative NLP models.
@@ 5.3.4 Alternative Terms for Hallucination
This indicates that while the majority of respondents did not propose an alternative term, a notable proportion sees fabrications as a more suitable descriptor for the phenomenon of hallucination in AI-generated content. For example,
@@ 5.3.5 Creativity and Positive Applications
Further supplementary analysis and quotes on the various external perspectives and the societal ramifications of hallucination, obtained through the survey, is examined in the Appendix 10.3.
6 Challenges and Recommendations
@@
Based on our audit and survey analysis, we outline the key weaknesses encountered in hallucination within NLP and potential recommendations motivated by the weaknesses. We utilize a community-centric approach to define the primary weaknesses of the field currently and a path forward.
@@ 6.1 Challenges
The primary challenges we identify thematically and aim to elucidate are as follows:
Wide range of vague and absent definitions: The literature and the practitioner’s survey show diverse and conflicting frameworks, often lacking clarity or omitting explicit definitions for hallucination and how it is perceived in various fields of NLP and language generation. Ambiguity arises from the use of terms like ‘confabulations,’ ‘ fabrications,’ ‘misinformation,’ and ‘hallucinations’ interchangeably, without clear definitions in the context of hallucinations.
Lack of standardization in measurement: The absence of standardized metrics to quantify hallucination results in the use of multiple scales and categorizations. This makes it challenging to compare and interpret results across different models and studies, leading to a proliferation of diverse approaches for evaluating hallucinations.
Limited awareness of hallucination in a sociotechnical context: Hallucination research often lacks the understanding of how the concept of hallucination is conceptualized beyond its technical purview. When such analysis is employed in sociotechnical systems like healthcare and policy making Dahl et al. (2024); Pal et al. (2023), it is important to define the socially relevant framework of hallucination as well. There is no motivation shown to understand the non-technical considerations of hallucination.
Multiple sentiment towards hallucination: The perception of hallucination in generative AI varies depending on the context. For instance, it is often positively regarded as creativity in image generation, whereas in text generation, it is viewed negatively as errors or mistakes. Consequently, future research efforts should aim to better address this disparity to develop a more nuanced framework for understanding hallucination.
@@
User trust and reliability: Our survey findings suggest that users may hesitate to fully utilize LLM capabilities due to concerns about bias and hallucination despite recognizing the potential advantages these models offer. Therefore, there is a need to focus efforts on understanding the human interaction aspect concerning hallucination in NLP and language generation.
@@ 6.2 Recommendations
Expanding on audits like Blodgett et al. (2020) & Venkit et al. (2023), we examine strategies for NLP practitioners studying ‘hallucination’ to overcome these challenges. We propose two overarching themes with four associated recommendations.
Author-Centric Recommendation.
@@
Community-Centric Recommendation.
@@ 7 Conclusion
Our work delves into the conceptualization of hallucination within the scope of NLP. Our approach involved two key methodologies: first, an exhaustive audit of 103 peer-reviewed papers in the NLP domain, and second, a practitioner survey of 171 researchers to complement our first study with real-world practical perception and understanding of hallucination as a unique contribution. Through this analysis, we have gained insights into how the NLP community conceptualizes and defines hallucination, showcasing a lack of discourse and agreement. Additionally, our thematic and community-based approach highlights potential weaknesses within the field, particularly in addressing misrepresentations and inaccurate characterizations associated with hallucination, paving way for better advancement in language generation.
@@ 9 Ethics Statement
Data Publication: All the papers used in our research are listed in the Appendix. However, we recognize the importance of transparency and accountability. Therefore, we publish the complete collection of papers along with our qualitative classification and annotation, allowing for public scrutiny and examination  [^2] https://github.com/PranavNV/The-Thing-Called-Hallucination.